The dividend on a common stock just paid this year was $2.25 while the dividend paid 10 years ago was $1.05. The current dividend payout ratio is 40% and the investor's required return is 10%. What is the current P/E ratio for this stock?

19.23
Answer #1

earnings = dividend/payout ratio

earnings = 2.25/0.4 = 5.625

growth in dividends = (2.25/1.05)^(1/10) - 1 = 0.07919350247

Price = Div1/(r - g)

Price = 2.25 * (1 + 0.07919350247) /(0.1 - 0.07919350247)

Price = 116.70322586

P/E = 116.70322586/5.625

P/E = 20.7472401529

= 20.75